a random sample of 150 worker with children in day care show a man day a care cost of r. s 3700 a standard deviation is r. s 600 verified the department claim that the mean is equal to r. s3500 at the 0.05 level of significance


1
Expert's answer
2021-02-24T06:52:41-0500

The provided sample mean is "\bar{x}=3700" and the sample standard deviation is "s=600."

The sample size is "n=150." The number of degrees of freedom are "df=n-1=149," and the significance level is "\alpha=0.05."

The following null and alternative hypotheses need to be tested:

"H_0: \mu=3500"

"H_1: \mu\not=3500"

This corresponds to two-tailed test, for which a t-test for one mean, with unknown population standard deviation will be used.

Based on the information provided, the critical value for a two-tailed test is "t_c=1.976013."

The rejection region for this two-tailed test is "R=\{t:|t|>1.976013\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{3700-3500}{600/\sqrt{150}}=4.082483"

Since it is observed that "t=4.082483>1.976013=t_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 3500, at the 0.05 significance level.

Using the P-value approach: The p-value "df=149, t=4.082483," "two-tailed," is "p=0.000072," and since "p=0.000072<0.05," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu" is different than 3500, at the 0.05 significance level.